The car with mass m = 1.2 t passes at a speed of 20 km / h through a bridge bent down. Calculate the force with which the compression machine bridges to the lowest point, when it is known that the radius of the bridge is R = 10 m.


1
Expert's answer
2021-01-16T17:24:46-0500

"mg-N=-m\frac{v^2}{r}\to N=mg+m\frac{v^2}{r}="


"=1200\cdot9.8+1200\cdot\frac{5.56^2}{10}=15464(N)" . Answer